RC meeting and Run Plan, Wednesday February 15th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Report
- https://logbooks.jlab.org/entry/4131947
- RC: Sean Jeffas, sj9ry@virginia.edu
Summary
- Ran with ~30 uA beam overnight increasing to 40 uA during the dat, collecting 0.8C
- Zeke and Anu got SBS GEMs working during the downtime - https://logbooks.jlab.org/entry/4131150
- SBS GEMs were included in the data overnight with minimal issues.
- At 2 pm one SBS GEM HV started acting up and is being investigated - https://logbooks.jlab.org/entry/4131946
Issues
- HCal ROC busy, simple reset fixed - https://logbooks.jlab.org/entry/4131218
- Hodoscope SL4 tripped twice, came back online right away - https://logbooks.jlab.org/entry/4131233
- One SBS GEM was tripping, needed to increase the trip time - https://logbooks.jlab.org/entry/4131581
- One GEM reset needed to fix the DAQ - https://logbooks.jlab.org/entry/4131631
- Currently one SBS GEM HV is no working - https://logbooks.jlab.org/entry/4131946
Current Status
- Running 45 uA on He3
Run Plan
- Continue running 45 uA on He3
- Accelerator will prioritize beam and not have any maintenance days this week
Data Collection and Polarization Status
- 31.8C of uncorrected charge for this kinematic (37.2C of corrected charge)
- Last NMR 54.6%%
Other Topics/Status Updates
- Anything else?
- Bill wants to change the raster settings after some analysis
Requested Accesses to the Hall
- Jack would like to collect some cables